OpenStudy (ttop0816):

The period of y = 2cotx is _____. π/2 π 2π <--- 4π

OpenStudy (amoodarya):

 so period = ?

OpenStudy (amoodarya):

\[f(x+T)=f(x) \rightarrow T = period\] \[2\cot(x+T)=2\cot(x) \rightarrow x+T=x+ k \pi so T=k \pi , the original period is when k=1\]

OpenStudy (ttop0816):

@amoodarya then it would be 2pi/1? equals to 2pi?

OpenStudy (amoodarya):

x+T=x+k pi T=k pi k=1 for original period so T=pi

OpenStudy (ttop0816):

the period is pi?